OpenStudy (anonymous):

How I find the range of the following, without graphing? \(\ \Huge \mathsf{\frac{x}{3x-1}} ? \)

OpenStudy (anonymous):

well the range is everywhere that y is defined.. Since this is a fraction the range is defined where the denominator is not 0.. so set 3x-1= 0 and solve for x

OpenStudy (anonymous):

So...

OpenStudy (anonymous):

That gives me the \(\ \mathsf{Domain}\) though, which is \(\ \Huge \mathsf{(-infin,\frac{1}{3})U(\frac{1}{3}, infin)}...\)

OpenStudy (anonymous):

you need to consider extreme values. consider this: \[\frac{x}{3x -1} = \frac{1}{3 - \frac{1}{x}}\] what happens when x has a very large magnitude?

OpenStudy (anonymous):

The 1/x gets closer to infity, so it's ~1/3

OpenStudy (anonymous):

does it ever equal 1/3 ?